2 2x +5x-3=0 ( i m yr new student my name is fareeha & i study in class 8 )....
Answer by stanbon(75887)

2. 2x +5x-3=0 ( i m yr new student my name is fareeha
& i study in class 8 )....
------------------
I'm going to assume your problem is the following:
2x+5x-3 = 0
---
Add the 2 terms with "x" to get:
7x - 3 = 0
Add 3 to both sides of the equation to get:
7x = 3
Divide both sides of the equation by 7 to get:
x = 3/7
=============
Cheers,
Stan H.